Let W be the set of vectors in R3 of the form(a,b,½a-2b)

show that W is subspace of R3

1
Expert's answer
2021-11-01T16:55:28-0400

Let m, n ∈ W and Let β∈ℝ. Then m = (a, b, ½a-2b) and n = (x, y, ½x-2y)


We shall show that m + βn ∈ W


m + βn = (a, b, ½a-2b) + β(x, y, ½x-2y)

= (a, b, ½a-2b) + (βx, βy, β(½x-2y))

= (a+βx, b+βy, ½a-2b + β(½x-2y))


Since a+βx, b+βy, ½a-2b + β(½x-2y) ∈ ℝ ∀ a, b, x, y, β ∈ ℝ, then m+βn ∈ ℝ.


Thus, W is a subspace of ℝ³